# Learning from galactic rotation curves: a neural network approach

### v1

@@ II.4 Pre-processing
$\tilde{{\bf v}}_{l}=\frac{{\bf v}_{l}-\mu_{l}}{\sigma_{l}}\ .$ (13)
Here $\mu_{l}$ and $\sigma_{l}$ are the mean and standard deviation over the $l$th radius value of the training data (validation and test data are not included in this calculation). The boldface here implies a column vector the size of the training data, i.e. $4\times 10^{5}$. We then use the mean and standard deviation of the training set itself to scale the validation and test sets, as well as the actual observed rotation curve before feeding them to the neural network.
@@
II.5 Neural network architecture
It is easy to see from the discussion in Section II.2.1 that a larger number of neurons per layer as well as a larger number of layers in a neural network will enable it to approximate a complex function better. However, the performance of the network also depends on other parameters like batch size, learning rate of the optimizer, choice of the optimization algorithm and activation function as well as the loss function used, etc. These parameters that characterize a neural network are called hyperparameters and choosing their optimal values for a neural network is a difficult task. This is because they must be chosen empirically, i.e. by training the neural network multiple times with various combinations of hyperparameters and choosing those which perform best on some unseen data.
@@ II.5 Neural network architecture
    Loss function: We use the mean-squared-error (MSE) loss given by Eq. (12). Note that we shall also implement a different loss function in Section IV.2 given by Eq. (15).
@@
    Batch size: For a stochastic gradient descent method, IAPs are updated based on loss calculated for a small subset of size $\mathcal{B}$ of the total training set. This small subset is called a batch (or a mini-batch), and is randomly drawn from the training set without replacement. We use $\mathcal{B}=32$, implying $125004\times 10^{5}/32=$ 12500 updates (or iterations) of the internal adjustable parameters before all samples from the training set are fed to the neural network once.
@@
    Dropout: To prevent overfitting we also utilize dropout regularization, where for every sample of the training set, the output of each neuron is set to zero with probability $d$, so that each sample is passed through a different ‘thinned’ neural network [46]. We set $d=0.2$ for every hidden layer in the network.
@@
To find the optimum architecture of the neural network, we first fix the above hyperparameters at the values mentioned. We then construct a grid that consists of the number of hidden layers $L\in[1,2,3,4]$ and neurons per layer $N_{j}\in[100,150,200]$. For a galaxy, for each combination of $L$ and $N_{j}$ we train a neural network for 250 epochs. The performance of these trained networks is then evaluated using the loss on validation data, (recall that the weights and biases are not updated based on validation data) and we choose the architecture for which the validation loss is the lowest. Note that this grid search is carried out with noise-induced simulated rotation curves (see Section III.2 for details) with the mean-squared-error loss function, while the obtained optimal architecture is utilized for the case of noiseless training data as well as the case with a heteroscedastic loss function (as we shall discuss in Section IV.2).
@@
We perform this grid search for $4$ galaxies, and find that validation loss is lowest for networks with $2$ or $3$ hidden layers with either $150$ or $200$ neurons per layer with very small differences in the loss between these architectures. We therefore fix the number of hidden layers to $2$ and neurons per layer to $200$ for all 7 galaxies in the sample.
@@
Finally, using the index notation discussed in Section II.2.1, the neural network architecture will comprise of an input layer ($j=0$), whose size is the number of observed radius values for a particular galaxy, $2$ hidden layers ($j=1,2$) with $n_{j}=200$ each and an output layer $j=3$ with $n_{j}=5$ outputs corresponding to a vector given by Eq. (7) in parameter space.
@@
III Inferring model parameters using mean-squared-error loss function
@@
Once the simulated rotation curves are obtained and the architecture is finalized, we can now train our neural networks. In this section, we train two neural networks for each galaxy in the sample: (a) without noise-induced simulated rotation curves and (b) noisy simulated rotation curves.
@@
For both neural networks, we use the mean-squared-error loss defined in Eq. (12) during training, while all other hyperparameters including the architecture are fixed to what was discussed in Section II.5.
III.1 Noiseless Case
@@
Let us start with the simpler case, where simulated rotation curves without any noise are used for training. We follow the pre-processing steps in Section II.4 and then train $7$ different neural networks, each for a galaxy in our sample of DM dominated dwarf galaxies. The neural networks are trained for $250$ epochs while both training and validation loss are monitored. We then test the performance of the network by calculating the loss on the test data. This is done to ensure that test loss is not too different from training loss, i.e. the neural network does just as well on unseen data as it does on training data.
@@
Since we do not have target parameters for the observed rotation curves to check how good these predictions are, we must utilize another metric to discern if the predicted parameters are good. One way to do so is to use the predicted parameter vector ${\bf\hat{P}}$ to construct a rotation curve using Eqs. (2) - (6), which can then be compared to the observed one using a $\chi^{2}_{red}$ (reduced $\chi^{2}$) value.
@@
  • To make sure performance on test data is good, along with calculating loss, one can also compare constructed rotation curves using the target parameters to those obtained from the predicted parameters. We do this for three randomly selected examples from the test data in Fig. 2. It can be seen that, while the rotation curves don’t match exactly (since the loss is not zero for parameter predictions), the predicted curves look visually similar and are close to the target rotation curves.
@@
  • However, when observed rotation curves are given as input, the predicted parameters are not able to reconstruct the observed rotation curve with a low $\chi^{2}_{red}$, which can be seen by the blue curve in Fig. 3. One reason for this could be that even without the error bars, central values of the observed velocities do not form a smooth rotation curve (see for instance, the dip in the observed velocity at $\sim 1\ \text{kpc}$ in Fig. 3). Since the dark matter component for simulated data is a smooth curve, these bumps are not learned by the neural network. This can be seen in other galaxies as well, where in Fig. 4, parameters for observed velocities (without uncertainties) which are smoother are predicted better by the corresponding neural network.
@@ III.2 Noisy case
In every simulated rotation curve, for each radius bin $r_{i}$, we add a random draw from the distribution $\mathcal{N}(0,\sigma(r_{i})^{2})$ as noise to the simulated velocity $v_{sim}(r_{i})$. Note that the target outputs, i.e. the known parameters, are not noisy and will not be affected by this process. The usual pre-processing steps are carried out as discussed in Section II.4.
@@
  • After training, we evaluate the performance of the neural network on test data, for which we know the target parameters. We find that the test loss is similar to the training and validation loss implying no overfitting. It is worth noting that due to the regularization effect of the input noise, the MSE loss for all data sets is higher compared to the noiseless case. However, the when rotation curves corresponding to target and predicted parameters are constructed, we find good agreement between the two as seen in Fig. 2.
  • On the other hand, unlike the noiseless case, we find that the performance on observed rotation curves has improved drastically. We feed as inputs, the central values of the observed velocities and construct rotation curves from the output parameters predicted. The corresponding curves and their $\chi^{2}_{red}$ are shown in Fig. 4 by the red curves. Note that the $\chi^{2}_{red}$ values are far smaller compared to the noiseless case, with predicted curves fitting much better than before.
Similar results are obtained for all the other galaxies in our sample, as seen in Fig. 4. In conclusion, neural networks trained with noisy training data perform much better those trained without noise. In the next section we shall discuss how one can also obtain the uncertainties associated with the parameters along with the point-estimates.
IV Estimating uncertainties in model parameters
@@
In this section, following recent work [11, 13, 15] we try to account for the aleatoric uncertainty in multiple ways, for neural networks trained on noisy inputs. First we explore how multiple realizations of the observed rotation curves can enable us to obtain multiple estimates of parameters in Section IV.1. These point-estimates can be used to construct ‘joint’ and ‘marginalized’ distributions for the predicted parameters. In Section IV.2 we also explore the use of a heteroscedastic loss function to implicitly learn the uncertainty during training itself.
IV.1 Using multiple realizations of the observations
@@
Thus, for every galaxy in our sample, we now have multiple realizations of the observed rotation curves for that galaxy. We can use these realizations as inputs to the trained neural network corresponding to each galaxy. Each realization will lead to a slightly different parameter prediction ${\bf P}$. We generate $1000$ realizations and obtain $1000$ predictions for each galaxy in the sample.
For every galaxy, we then consider the $50\%$ quantile (i.e., median) for each parameter as the point-estimate and $16\%$, $84\%$ quantiles (corresponding to a $1\sigma$ region of a $1$D normal distribution) as the uncertainties. This definition will help us in comparing our results to those obatined using MCMC in Section V. Note that a similar procedure was used in [11] to estimate cosmological parameters from CMB observations. The resultant parameters and their uncertainties are shown in Table 2. We also plot the rotation curves constructed from predictions and obtain the goodness of fit to the observations using $\chi^{2}_{red}$ (reduced $\chi^{2}$). The rotation curves are shown in Fig. 7 by the orange curves. It is clear that similar to the case with a single realization of the observed rotation curve, the rotation curves corresponding to the mean value of parameters fit observations just as well, or maybe even better, especially in the case of DDO 154 and UGCA 444.
The method of multiple realizations that we have discussed here gives us a sequence of parameters which can then be used to construct joint and marginal distributions. One could then compare the posteriors obtained from a likelihood-based MCMC sampling approach. In Section V we do this for UGC 5721 and compare the contours with those obtained using the approach described in this section.
Galaxy Predictions
@@
DDO 154 $1.96_{-0.16}^{+0.12}$ $5380.39_{-103.57}^{+101.64}$ $3.63_{-0.69}^{+0.85}$ $6.37_{-0.57}^{+0.63}$ $0.61_{-0.08}^{+0.05}$
ESO 444-G084 $5.06_{-0.98}^{+0.69}$ $4954.22_{-157.20}^{+123.29}$ $1.26_{-0.08}^{+0.2}$ $9.91_{-1.06}^{+0.72}$ $0.58_{-0.01}^{+0.01}$
UGC 5721 $2.28_{-0.30}^{+0.37}$ $3042.17_{-127.19}^{+155.87}$ $2.44_{-0.34}^{+0.57}$ $7.74_{-1.76}^{+0.94}$ $0.65_{-0.04}^{+0.03}$
UGC 5764 $3.97_{-0.42}^{+0.35}$ $4658.97_{-119.27}^{+122.23}$ $1.92_{-0.32}^{+0.64}$ $4.18_{-1.00}^{+1.83}$ $0.56_{-0.01}^{+0.01}$
UGC 7524 $1.90_{-0.43}^{+0.62}$ $4403.68_{-256.26}^{+376.14}$ $3.02_{-0.75}^{+1.30}$ $10.35_{-1.7}^{+1.29}$ $0.57_{-0.03}^{+0.03}$
UGC 7603 $2.59_{-0.27}^{+0.47}$ $4128.01_{-177.46}^{+326.71}$ $2.29_{-0.42}^{+0.47}$ $7.35_{-0.36}^{+0.68}$ $0.54_{-0.04}^{+0.05}$
UGCA 444 $7.96_{-0.95}^{+1.05}$ $7178.33_{-510.66}^{+384.62}$ $1.74_{-0.01}^{+0.01}$ $7.93_{-0.02}^{+0.02}$ $0.55_{-0.001}^{+0.001}$
Table 2: Point-estimates (median) and uncertainties ($1\sigma$ region) of $1000$ inferred parameters obtained from $1000$ realizations of the observed rotation curves for each galaxy.
IV.2 Employing heteroscedastic loss function

### v2

@@ 2.4 Pre-processing
$\tilde{{\bf v}}_{l}=\frac{{\bf v}_{l}-\mu_{l}}{\sigma_{l}}\ .$ (2.13)
Here $\mu_{l}$ and $\sigma_{l}$ are the mean and standard deviation over the $l$th radius value of the training data (validation and test data are not included in this calculation). The boldface here implies a column vector the size of the training data, i.e. $4\times 10^{5}$. We then use the mean and standard deviation of the training set itself to scale the validation and test sets, as well as the observed rotation curve before feeding them to the neural network.
@@
2.5 Neural network architecture
It is easy to see from the discussion in section 2.2.1 that a larger number of neurons per layer as well as a larger number of layers in a neural network will enable it to approximate a complex function better. However, the performance of the network also depends on other parameters like batch size, learning rate of the optimizer, choice of the optimization algorithm and activation function as well as the loss function used, etc. These parameters that characterize a neural network are called hyperparameters and choosing their optimal values for a neural network is a difficult task. This is because they must be chosen empirically, i.e. by training the neural network multiple times with various combinations of hyperparameters and choosing those which perform best on unseen data.
@@ 2.5 Neural network architecture
    Loss function: We use the mean-squared-error (MSE) loss given by eq. (2.12). Note that we shall also implement a different loss function in section 4.2 given by eq. (4.1).
@@
    Batch size: For a stochastic gradient descent method, IAPs are updated based on loss calculated for a small subset of size $\mathcal{B}$ of the total training set. This small subset is called a batch (or a mini-batch), and is randomly drawn from the training set without replacement. We use $\mathcal{B}=32$, implying $4\times 10^{5}/32=12500$ updates (or iterations) of the internal adjustable parameters before all samples from the training set are fed to the neural network once.
@@
    Dropout: To prevent overfitting we also utilize dropout regularization, where for every sample from the training set, the output of each neuron is set to zero with probability $d$, so that each sample is passed through a different ‘thinned’ neural network [52]. We set $d=0.2$ for every hidden layer in the network.
@@
To find the optimum architecture of the neural network, we first fix the above hyperparameters at the values mentioned. We then construct a grid that consists of the number of hidden layers $L\in[1,2,3,4]$ and neurons per layer $N_{j}\in[100,150,200]$. For a galaxy, for each combination of $L$ and $N_{j}$ we train a neural network for 250 epochs. The performance of these trained networks is then evaluated using the loss on validation data, (recall that the weights and biases are not updated based on validation data) and we choose the architecture for which the validation loss is the lowest. Note that this grid search is carried out with noise-induced simulated rotation curves (see section 3.2 for details) with the mean-squared-error loss function, while the obtained optimal architecture is also utilized for the case of noiseless training data as well as the case with a heteroscedastic loss function (as we shall discuss in section 4.2).
@@
We perform this grid search for $4$ galaxies, and find that validation loss was lowest for networks with at least $2$ hidden layers, while the number of neurons per layer varied. It is worth noting that the difference between validation loss for the best performing architectures for a given galaxy was very small ($\mathcal{O}(10^{-3})$). Therefore for purposes of this paper, we fix the number of hidden layers to $2$ and neurons per layer to $200$ for all 7 galaxies in the sample.
@@
Finally, using the index notation discussed in section 2.2.1, the neural network architecture will comprise of an input layer ($j=0$), whose size is the number of observed radius values for a particular galaxy, $2$ hidden layers ($j=1,2$) with $N_{j}=200$ each and an output layer $j=3$ with $N_{3}=5$ outputs corresponding to a vector given by eq. (2.7) in parameter space.
@@
3 Inferring model parameters using mean-squared-error loss function
@@
Once the simulated rotation curves are obtained and the architecture is finalized, we can now train our neural networks. In this section, we train two neural networks for each galaxy in the sample: one without noise included in the inputs of the training set (i.e. simulated rotation curves) and one with noise included.
@@
For both neural networks, we use the mean-squared-error loss defined in eq. (2.12) during training, while all other hyperparameters including the architecture are fixed to what was discussed in section 2.5.
3.1 Noiseless Case
@@
Let us start with the simpler case, where simulated rotation curves without any noise are used for training. We follow the pre-processing steps in section 2.4 and then train $7$ different neural networks, each for a galaxy in our sample of DM dominated dwarf galaxies. The neural networks are trained for $250$ epochs while both training and validation loss are monitored. The loss as a function of epochs is plotted in Appendix A in Figs. 10 and 11 for all galaxies. We then test the performance of the network by calculating the loss on the test data. This is done to ensure that test loss is not too different from training loss, i.e. the neural network does just as well on unseen data as it does on training data.
@@
Since we do not have target parameters for the observed rotation curves to check how good these predictions are, we must utilize another metric to discern if the predicted parameters are good. One way to do so is to use the predicted parameter vector $^{\bf\hat{P}}$ to construct a rotation curve using eqs. (2.2) - (2.6), which can then be compared to the observed one using a $\chi^{2}_{red}$ (reduced $\chi^{2}$) value.
@@
  • To make sure performance on test data is good, along with calculating loss, one can also compare constructed rotation curves using the target parameters to those obtained from the predicted parameters. We do this for three randomly selected examples from the test data in figure 2. It can be seen that, while the rotation curves don’t match exactly (since the loss is not zero for parameter predictions), the predicted curves look visually similar and are close to the target rotation curves.
@@
  • However, when observed rotation curves are given as input, the predicted parameters are not able to reconstruct the observed rotation curve with a low $\chi^{2}_{red}$, which can be seen by the blue curve in figure 3. One reason for this could be that even without the error bars, central values of the observed velocities do not form a smooth rotation curve (see for instance, the dip in the observed velocity at $\sim 1\ \text{kpc}$ in figure 3). Since the dark matter component for simulated data is a smooth curve, these bumps are not learned by the neural network. This can be seen in other galaxies as well, where in figure 4, parameters for observed velocities (without uncertainties) which are smoother are predicted better by the corresponding neural network.
@@ 3.2 Noisy case
In every simulated rotation curve, for each radius bin $r_{i}$, we add a random draw from the distribution $\mathcal{N}(0,\sigma(r_{i})^{2})$ as noise to the simulated velocity $v_{sim}(r_{i})$. Note that the target outputs, i.e. the known parameters, are not noisy and will not be affected by this process. The usual pre-processing steps are carried out as discussed in section 2.4.
@@
  • After training, we evaluate the performance of the neural network on test data, for which we know the target parameters. We find that the test loss is similar to the training and validation loss implying no overfitting. It is worth noting that due to the regularization effect of the input noise, the MSE loss for all data sets is higher compared to the noiseless case. However, when rotation curves corresponding to target and predicted parameters are constructed, we find good agreement between the two as seen in figure 2.
  • On the other hand, unlike the noiseless case, we find that the performance on the observed rotation curve has improved drastically. We feed as input, the central values of the observed velocities and construct the rotation curve from the output parameters predicted. The corresponding curve and $\chi^{2}_{red}$ are shown in figure 3 in red. Note that the $\chi^{2}_{red}$ value is far smaller compared to the noiseless case.
Similar results are obtained for all the other galaxies in our sample, as seen in figure 4. In conclusion, neural networks trained with noisy training data perform much better than those trained without noise. In the next section we shall discuss how one can also obtain the uncertainties associated with the parameters along with the point-estimates.
4 Estimating uncertainties in model parameters
@@
In this section, following recent work [11, 13, 15] we try to account for the aleatoric uncertainty in multiple ways, for neural networks trained on noisy inputs. First we explore how multiple realizations of the observed rotation curves can enable us to obtain multiple estimates of parameters in section 4.1. These point-estimates can be used to construct ‘joint’ and ‘marginalized’ distributions for the predicted parameters. In section 4.2 we also explore the use of a heteroscedastic loss function to implicitly learn the uncertainty during training itself.
4.1 Using multiple realizations of the observations
@@
Thus, for every galaxy in our sample, we now have multiple realizations of the observed rotation curves for that galaxy. We can use these realizations as inputs to the trained neural network corresponding to each galaxy. Each realization will lead to a slightly different parameter prediction $^{\bf\hat{P}}$. We generate $1000$ realizations and obtain $1000$ predictions for each galaxy in the sample.
For every galaxy, we then consider the $50\%$ quantile (i.e., median) for each parameter as the point-estimate and $16\%$, $84\%$ quantiles (corresponding to a $1\sigma$ region of a 11D normal distribution) as the uncertainties. This definition will help us in comparing our results to those obatined using MCMC in section 5. Note that a similar procedure was used in [11] to estimate cosmological parameters from CMB observations. The resultant parameters and their uncertainties are shown in Table 2. We also plot the rotation curves constructed from predictions and obtain the goodness of fit to the observations using $\chi^{2}_{red}$ (reduced $\chi^{2}$). The rotation curves are shown in figure 7 by the purple dashed curves. It is clear that similar to the case with a single realization of the observed rotation curve, the rotation curves corresponding to the median value of parameters fit observations just as well, or even better, especially in the case of DDO 154 and UGCA 444.
Galaxy Method Predictions
@@
DDO 154 MSE $1.96_{-0.16}^{+0.12}$ $5380.39_{-103.57}^{+101.64}$ $3.63_{-0.69}^{+0.85}$ $6.37_{-0.57}^{+0.63}$ $0.61_{-0.08}^{+0.05}$
Hetero $1.95_{-0.30}^{+0.30}$ $5350.71_{-292.16}^{+292.16}$ $3.50_{-0.87}^{+0.87}$ $6.87_{-3.90}^{+3.90}$ $0.61_{-0.12}^{+0.12}$
MCMC $1.79_{-0.10}^{+0.16}$ $5328.85_{-90.48}^{+136.65}$ $3.25_{-0.83}^{+1.61}$ $4.01_{-1.56}^{+3.56}$ $0.70_{-0.14}^{+0.08}$
ESO 444-G084 MSE $5.06_{-0.98}^{+0.69}$ $4954.22_{-157.20}^{+123.29}$ $1.26_{-0.08}^{+0.2}$ $9.91_{-1.06}^{+0.72}$ $0.58_{-0.01}^{+0.01}$
Hetero $5.26_{-1.07}^{+1.07}$ $5055.20_{-363.85}^{+363.85}$ $1.26_{-0.27}^{+0.27}$ $9.53_{-3.38}^{+3.38}$ $0.56_{-0.14}^{+0.14}$
MCMC $5.29_{-1.02}^{+0.94}$ $5181.16_{-324.19}^{+257.23}$ $1.14_{-0.11}^{+0.21}$ $10.39_{-3.63}^{+3.15}$ $0.58_{-0.18}^{+0.16}$
UGC 5721 MSE $2.28_{-0.30}^{+0.37}$ $3042.17_{-127.19}^{+155.87}$ $2.44_{-0.34}^{+0.57}$ $7.74_{-1.76}^{+0.94}$ $0.65_{-0.04}^{+0.03}$
Hetero $2.21_{-0.34}^{+0.34}$ $3052.34_{-163.26}^{+163.26}$ $2.55_{-0.72}^{+0.72}$ $7.12_{-4.04}^{+4.04}$ $0.66_{-0.13}^{+0.13}$
MCMC $2.12_{-0.22}^{+0.30}$ $3064.23_{-99.85}^{+124.44}$ $2.56_{-0.41}^{+0.36}$ $7.27_{-4.74}^{+5.20}$ $0.70_{-0.13}^{+0.07}$
UGC 5764 MSE $3.97_{-0.42}^{+0.35}$ $4658.97_{-119.27}^{+122.23}$ $1.92_{-0.32}^{+0.64}$ $4.18_{-1.00}^{+1.83}$ $0.56_{-0.01}^{+0.01}$
Hetero $3.87_{-0.43}^{+0.43}$ $4769.11_{-192.48}^{+192.48}$ $2.00_{-0.51}^{+0.51}$ $4.75_{-3.74}^{+3.74}$ $0.57_{-0.14}^{+0.14}$
MCMC $4.25_{-0.35}^{+0.43}$ $4879.48_{-106.92}^{+123.52}$ $1.40_{-0.30}^{+0.28}$ $1.42_{-0.32}^{+0.90}$ $0.56_{-0.18}^{+0.17}$
UGC 7524 MSE $1.90_{-0.43}^{+0.62}$ $4403.68_{-256.26}^{+376.14}$ $3.02_{-0.75}^{+1.30}$ $10.35_{-1.70}^{+1.29}$ $0.57_{-0.03}^{+0.03}$
Hetero $1.55_{-0.52}^{+0.52}$ $4370.11_{-422.96}^{+422.96}$ $3.02_{-1.43}^{+1.43}$ $9.56_{-2.86}^{+2.86}$ $0.57_{-0.14}^{+0.14}$
MCMC $1.36_{-0.28}^{+0.70}$ $4196.24_{-315.63}^{+609.70}$ $2.48_{-0.89}^{+1.22}$ $10.35_{-3.22}^{+2.92}$ $0.64_{-0.17}^{+0.12}$
UGC 7603 MSE $2.59_{-0.27}^{+0.47}$ $4128.01_{-177.46}^{+326.71}$ $2.29_{-0.42}^{+0.47}$ $7.35_{-0.36}^{+0.68}$ $0.54_{-0.04}^{+0.05}$
Hetero $2.61_{-0.65}^{+0.65}$ $4243.85_{-368.33}^{+368.33}$ $2.20_{-0.77}^{+0.77}$ $7.43_{-4.04}^{+4.04}$ $0.56_{-0.14}^{+0.14}$
MCMC $2.48_{-0.42}^{+0.62}$ $4186.56_{-221.71}^{+345.33}$ $2.18_{-0.83}^{+0.74}$ $6.03_{-3.56}^{+5.78}$ $0.57_{-0.17}^{+0.15}$
UGCA 444 MSE $7.96_{-0.95}^{+1.05}$ $7178.33_{-510.66}^{+384.62}$ $1.74_{-0.01}^{+0.01}$ $7.93_{-0.02}^{+0.02}$ $0.55_{-0.001}^{+0.001}$
Hetero $6.84_{-1.61}^{+1.61}$ $7958.19_{-742.22}^{+742.22}$ $1.47_{-0.35}^{+0.35}$ $8.23_{-3.91}^{+3.91}$ $0.57_{-0.14}^{+0.14}$
MCMC $6.90_{-1.40}^{+1.13}$ $8467.32_{-640.01}^{+386.14}$ $1.34_{-0.22}^{+0.47}$ $8.57_{-4.56}^{+4.38}$ $0.57_{-0.18}^{+0.16}$
Table 2: Parameters and their uncertainties obtained by feeding observed rotation curves as input to neural networks trained using: (a) the multiple realization method using MSE loss function (section 4.1), (b) a heteroscedastic loss function (section 4.2) and (c) using MCMC (section 5).
The method of multiple realizations that we have discussed here gives us a sequence of parameters which can then be used to construct joint and marginal distributions. One could then compare these distributions to the posteriors obtained from a likelihood-based MCMC sampling approach. In section 5 we do this for UGC 5721 and compare the contours with those obtained using the approach described in this section.
4.2 Employing heteroscedastic loss function